### v1

# The Winner of the NFL Draft is Not Necessarily Cursed

Abstract
Football analysts traditionally determine the relative value of draft picks by average future player value at each draft position. One implication is the loser’s curse: top draft picks belonging to last year’s worst teams produce less surplus value on average than draft picks later in the first round belonging to better teams. Additionally, these valuations do not match the valuation implied by the trade market. Either general managers are making terrible trades on average, or there is a sound economic reason for the discrepancy; we are partial to the latter explanation. Traditional analyses don’t consider that variance in performance decays convexly accross the draft, causing eliteness (e.g., right tail probability) to decay much more steeply than expected value. Because elite players have an outsize influence on winning the Super Bowl, we suspect general managers value performance nonlinearly, placing exponentially higher value on players as their eliteness increases. Draft curves that account for this closely resemble the trade market. Additionally, we create draft curves that adjust for position via a novel Bayesian hierarchical Beta regression model. We find that if you are interested in an elite quarterback, there is no loser’s curse.
Keywords: Decision making under uncertainty, applications and case studies, statistics in sports, NFL Draft, Beta regression, Bayesian hierarchical model
1 Introduction
1.1 Background information
The NFL Draft. Players joining the National Football League (NFL) are allocated to teams via an annual draft. Teams take turns selecting players in an order determined by their win-loss records from the previous year. In each of seven rounds, teams select players in increasing order of the previous year’s win-loss record. Drafting begins with last year’s worst team and ends with last year’s champion (with some minor exceptions). Each drafted player is then signed to a four-year first contract, also known as the rookie contract. Teams can choose to add a fifth year to the contracts of first-round selections, known as a fifth-year team option.
Player quality decays convexly over the course of the draft. Players drafted in the first round often turn out to be solid or excellent NFL players. Players drafted in the seventh round rarely turn out to be solid NFL players. Reflecting this trend, a drafted player’s first contract salary is determined by his draft position. First contract compensation decreases convexly over the first few rounds of the draft, leveling off after the fourth round. First contract compensation is essentially deterministic, although drafted rookies sometimes negotiate elements of the contract such as performance bonuses or injury guarantees. These rookie contracts count towards the NFL salary cap, a rule that determines the maximum amount any team can pay its players in a given year. Over time the cap has increased from \$123 million in 2013 to \$224.8 million in 2023. But, rookies are generally paid much less than veteran players who sign in free agency. In free agency, NFL players who are not bound by a contract can sign a new contract with any team.
The Market for NFL Draft Picks. Trading draft picks is a big part of football. NFL teams often find themselves wanting to trade draft picks, and such trades typically consist of current picks, future picks, and/or active players. A general manager may want to trade up to draft a particular player before he is taken by another team. For example, in 2017 the Kansas City Chiefs sent a first-round pick (the $27^{th}$ pick), a third-round pick (the $91^{st}$ pick), and its 2018 first-round pick to the Buffalo Bills in exchange for the $10^{th}$ pick, which it used just minutes after the trade to draft quarterback Patrick Mahomes. A general manager may also want to trade up to be able to draft players of a certain caliber at a particular position. For example, in 2023 the Carolina Panthers sent a first-round pick (the $9^{th}$ pick), a second-round pick (the $61^{st}$ pick), its 2024 first-round pick, its 2025 second-round pick, and wide receiver DJ Moore to the Chicago Bears for the $1^{st}$ overall pick, with the intention of drafting an elite quarterback.
In 2017, how did the Chiefs and the Bills decide that the draft-day trade was worth executing? More generally, if a team wants to trade for a particular draft pick, which picks should it offer in exchange? And if a team is offered a bundle of draft picks in exchange for another bundle of draft picks, should it accept or reject the trade? We are interested in the relative value of draft picks.
1.2 Traditional draft value curves
Football analysts who calculate the relative value of draft picks typically display their valuations as charts or curves, plotting value against draft position (e.g., see Figure 1 later in this section).
Jimmy Johnson’s Chart. In the 1990s, at the request of Dallas Cowboys coach Jimmy Johnson, vice president Mike McCoy created the first draft value curve based on gut instinct and looking at past trades. Known today as the Jimmy Johnson chart,111 https://www.the33rdteam.com/jimmy-johnsons-draft-trade-value-chart-set-standard-but-clearly-outdated/ it assigns a numerical point value to each draft position $1,...,256$. The first pick is worth the most, and the curve decays steeply and convexly. Jimmy Johnson’s formulation, and draft curves today, are additive: the value of a bundle of picks is the sum of the values of each individual pick. A trade is considered fair if the sum of the values of the two sides of the trade is equal. For example, consider the 2017 draft trade between the Chiefs and Bills, discussed earlier. According to Jimmy Johnson’s chart, the Bills gained value by accepting the trade: the Chiefs three picks are worth about 1,816 points,222 The Chiefs three picks are worth 680, 136, and 1000 points, respectively, supposing the Chiefs’ 2018 first-round pick is in the middle of the first-round at pick 16, which is reasonable considering the Chiefs made the playoffs 3 of the 4 years before 2017. more than the 1,300 point value of the Bills’ $10^{th}$ pick.
Though Jimmy Johnson’s chart is ad hoc and nonrigorous, general managers used it for decades, and some teams still use it today. More recently, Massey and Thaler, (2013) created rigorous data-driven draft value curves. In particular, they considered the expected value of a drafted player’s performance, and performance relative to cost, as a function of draft position.
Measuring player performance. Defining the value of a player’s performance in the first place takes care and consideration. Each play in American football consists of a complex interplay between 22 players of varying roles interacting in continuous time and space. This makes it difficult to measure how a player’s individual performance impacts the ultimate outcome of a play. For instance, a linebacker may impact a play simply by occupying space. He could cause the ball carrier to move in a different direction, and there is no easy way to measure the impact of that action. There is no easy or industry-standard way of measuring player performance on a given play in American football, particularly in a way that measures players of all positions on the same scale. Hence, analysts use performance proxies. Each measure is imperfect and has its own drawbacks. We describe many of these performance measures in detail in Appendix B. Throughout this paper, for simplicity, we use one performance outcome variable, detailed below.
In free agency, NFL players who are not bound by a contract can sign a new contract with any team. We believe the free agent market is relatively efficient and thus reflects the performance value of players. Therefore, following Baldwin, (2024), in this study we measure performance value by the average salary per year of a contract signed in free agency. We consider salary as a percentage of the cap, known as APY cap percentage, to adjust for inflation.
Football analysts, beginning with Massey and Thaler, (2013), have been interested in particular in a player’s first contract performance value rather than his career performance value. This is because an NFL player is paid particularly cheaply during his first contract. A rookie’s first contract is determined by his draft position and can be much lower than the value his performance provides, creating a first contract “surplus value.” Teams can take advantage of this extra cap space by signing good veteran players in free agency. The drawback of focusing solely on a player’s first contract is that the very best players stay with the same team across multiple contracts, providing value across their entire careers (e.g., Tom Brady). Continuing with the precedent set by Massey and Thaler, (2013) and subsequent analyses, in this study we consider just first contract performance and surplus value.
Thus, throughout this paper, we measure a player’s first contract performance by his second contract’s average salary per year as a percentage of the cap (APY cap percentage). This is a reflection of a player’s first contract performance value since he signs his second contract in free agency just after his first contract ends. Players who don’t sign a second contract have 0 performance value. We acknolwedge that free agent APY cap percentage is an imperfect measure of a player’s performance (see Appendix B) and first contract value is an incomplete measure of the total value he provides to his team (as it excludes the rest of his career). Nonetheless, the focus of this paper is to value draft picks given a performance measure, not to choose the best performance measure. We leave further analysis of various measures of performance to future work.
Traditional expected value curves. Now, we discuss the methodology underlying traditional data-driven draft curves. Formally, prior to the draft, we do not know how well a player drafted at pick $x\in\{1,...,256\}$ will perform in the NFL. Thus, we think of the performance outcome $Y$ associated with pick $x$ as a random variable, denoted by a capital letter. Recall that, throughout this paper, we let the realization of $Y$ be a player’s annualized second contract value relative to the salary cap (APY cap percentage). Then, from the dataset $\{(x_{i},Y_{i})\}$ of all recent draft picks in NFL history, we want to learn the de-noised relationship between pick number $x$ and performance outcome $Y$. Traditional analyses use a standard approach but with different measures of performance, detailed in Appendix B (Massey and Thaler,, 2013; Stuart,, 2012; Fitzgerald and Spielberger,, 2024; Pro Football Focus,, 2024; Baldwin,, 2024). The de-noised relationship they estimate from data is the expected value of $Y$ given $x$.
This traditional approach proceeds as follows. First, fit the expected performance curve, or the conditional mean $x\mapsto{\mathbb{E}}[Y|x]$. Second, calculate the cost curve, $x\mapsto{\mathrm{cost}}(x)$. Cost is first contract compensation, which is essentially a deterministic function of $x$. Third, calculate the expected surplus curve, or the expected value of the difference between the performance and cost, $x\mapsto{\mathbb{E}}[Y-{\mathrm{cost}}(x)|x]$. Finally, normalize each of these curves so that the value of the first pick is 1. Analysts typically normalize by dividing by the value of the first pick.
We fit these traditional draft value curves using our outcome variable $Y$ (second contract APY cap percentage), which we visualize in Figure 1. We use spline regression to fit the expected performance curve from a dataset of all draft picks from 2013 to 2023 (Ho and Carl,, 2024). We calculate the cost curve as in Baldwin, (2024): beginning with the actual 2023 salary cap (about \$$\$225$ million) and assuming a cap growth rate of 7% over the next three seasons (2024, 2025, and 2026), we convert the total dollar amount of the rookie contract into a percentage of the cap. Then, for each draft pick we average the compensation values across the four seasons. We calculate the expected surplus curve as the difference between the expected performance curve and cost curve.
Figure 1: Each curve is a graph of the value $v(x)$ of draft position $x$ relative to the first pick. In blue is the expected performance curve, proportional to $x\mapsto{\mathbb{E}}[Y|x]$. In red is the cost curve, proportional to $x\mapsto{\mathrm{cost}}(x)$. In green in the expected surplus curve, proportional to $x\mapsto{\mathbb{E}}[Y-{\mathrm{cost}}(x)|x]$. In orange is the draft value curve implied by the trade market. In pink is Jimmy Johnson’s curve.
Expected performance value (blue) is less steep than cost (red), indicating that players are underpaid on average in their first contracts. This produces positive expected surplus value (green) across most of the draft. Expected surplus value (green) is larger than 1 throughout the first round, peaking in the middle of the first round. This surplus spike is the “loser’s curse” from Massey and Thaler, (2013): top draft picks, belonging to last year’s worst teams, produce less surplus value on average than draft picks later in the first round that belong to better teams.
We also visualize Jimmy Johnson’s draft curve (pink) and the draft curve fit from the trade market (orange). Devised in Massey and Thaler, (2013), the fitted trade market curve is a singular draft value curve that best fits the dataset of all observed trades. We derive this curve in detail in Appendix C.
The fitted trade market and Jimmy Johnson curves are quite similar. This is not a coincidence: even today, some teams still trade using the Jimmy Johnson chart. The trade market curve is much steeper than expected performance and surplus value curves. In other words, teams aren’t valuing draft picks by their expected value. The implication is that general managers are trading suboptimally: they don’t trade down often enough, and they often overpay when they trade up. According to surplus value, even teams at the very top of the draft should trade down, particularly towards the middle of the first round. This result is counterintuitive because very top picks are so highly regarded by general managers and football fans.
1.3 Research hypothesis
Massey and Thaler, (2013) pioneered the traditional analytical approach to calculating NFL draft curves, which are expected value curves. They had two main findings. They discovered the loser’s curse: according to expected surplus value, top picks aren’t as valuable as later picks. Also, the way general managers trade indicates that they don’t value picks by expected performance or surplus value. In particular, general managers trade suboptimally: trading down in the draft can be incredibly valuable, and they often overpay when they trade up.
As shown in Section 1.2, Massey and Thaler’s analysis from over a decade ago completely replicates to recent data (2013-2023). Even though their work has been public and widely known for a long time, general managers’ draft behavior has remained largely unchanged. One explanation is that general managers are stubborn and haven’t learned from these analyses. If you believe general managers should value picks by expected performance or surplus value, they have continued to make terrible trades on average, not trading down often enough and overpaying when they trade up. An alternative explanation is that there is a sound economic reason why average future player value (at a draft position) doesn’t determine trade value. General managers are using some other valuation function that captures an essential piece of the puzzle missing from previous analyses. We are partial to the latter explanation.
A general manager’s goal is to win the Super Bowl. Winning the Super Bowl entails being the best team of 32. Mathematically, thinking of each team’s performance as a random variable, being the best team is like exceeding a high threshold of team performance. Having the highest expected value is not necessarily the best way to accomplish this. If high expected value comes with little variance, you may never exceed a high threshold.
Consider a toy example where the performance of each football team is determined by a Guassian random variable. Suppose team $A$’s performance has mean 1 and s.d. 0 and team $B$’s has mean 0 and s.d. 1. Suppose also that to have a good shot at winning the Super Bowl, a team needs to exceed a performance value of $1.28$. Though team $A$’s performance is much higher than $B$’s on average, $A$’s probability of exceeding the threshold is 0 and $B$’s is 10%. Team $B$ has a decent shot because its high variance puts nontrivial probability mass on far-right tail performance.
With this in mind, we consider how variance of performance changes as the draft progresses. In Figure 2 we visualize the conditional mean $x\mapsto{\mathbb{E}}[Y|x]$ and standard deviation $x\mapsto{\mathrm{sd}}(Y|x)$ of performance $Y$ as a function of draft position $x$. These graphs look remarkably similar. The key insight is that the conditional mean and variance of performance $Y$ decreases as draft pick $x$ increases. Moreover, they decrease convexly. Since an earlier pick has a much higher expected value and variance than a later pick, it has a much fatter right tail. In other words, an earlier pick has a significantly higher probability of becoming an elite player, moreso than we would expect if we just considered expected value.
Figure 2: The black dots indicate the empirical mean (left) and s.d. (right) of $Y$ (performance, i.e. second contract APY cap percentage) given $x$ (draft position). The blue lines are smoothed curves $x\mapsto{\mathbb{E}}[Y|x]$ (left) and $x\mapsto{\mathrm{sd}}(Y|x)$ (right).
If a general manager’s goal is to win a Super Bowl, then accepting a trade because it yields higher performance or surplus value on average is not necessarily optimal. We suspect general managers also value variance, which puts subtantial probability mass on far-right tail outcomes. In other words, we suspect general managers value performance nonlinearly, placing exponentially higher value on players as their eliteness increases. Intuitively, elite players in the far-right tail of performance are essential to a team’s performance exceeding a high threshold.
Ideally we would quantify the probability of winning a Super Bowl if a team drafts a player of performance value $Y$, but this is extremely complex for many reasons and we leave it to future work (see Section 3 for details). Instead, in this paper we consider simpler alternative value functions beyond expected performance or surplus value. We focus on nonlinear transformations of $Y$ that capture the outsize influence of elite players. The most intuitive example is right tail probability, which places high value on elite performance and no value on lesser performance. We find that such nonlinear value functions produce steeper draft value curves that more closely resemble the observed trade market. If you are looking for elite talent, picks from lower draft picks aren’t worth much and the high cost of a top draft pick is justified. Additionally, we adjust for position. In particular, we build draft value curves for quarterbacks and non-quarterbacks. We find that, if you are looking for an elite quarterback, there is no loser’s curse. The far-right tails of performance $Y$ are so fat for top pick quarterbacks that even the high cost of top picks doesn’t massively diminish their value.
The remainder of this paper is organized as follows. In Section 2 we build draft value curves that value eliteness and variance. We begin with right tail probability and extend to more intricate valuations. We then consider surplus value and adjust for position. In Section 3 we conclude.
2 Draft curves that value eliteness
2.1 Right tail probability
An intuitive example of valuing a draft position by eliteness is right tail probability. Formally, we consider the probability that the performance $Y$ of a player drafted at position $x$ exceeds a threshold $r$, ${\mathbb{P}}(Y>r|x)$. The tail cutoff $r$ is defined by a percentage of the salary cap, not a quantile. For example, a general manager may be interested in the probability that a drafted player is at least as valuable as Cowboys’ quarterback Dak Prescott ($r=15.8\%$). Since the expected value of an indicator is its probability, this is equivalent to the number of elite players expected to arise from a given pick.
We compute right tail probability from the conditional density ${\mathbb{P}}(Y|x)$, which we estimate using a Bayesian Beta regression model, detailed in Appendix D.1. As both the mean and variance of $Y$ decay convexly over the draft (recall Figure 2), the right tail decays convexly as well, which we visualize in Figure 3. The densities of top picks feature fat right tails, which shift lefward and morph into a spike near 0 as the draft progresses (see Figures 12 and 14 in Appendix D.1).
Figure 3: Right tail probability ${\mathbb{P}}(Y>r|x)$ ($y$-axis) as a function of draft position $x$ and eliteness cutoff $r$ (color).
In Figure 4 we visualize right tail probability draft value curves. We overlay the expected performance value curve (from Figure 1) and the fitted trade market curve (from Appendix C). The right tail curves are steeper than the expected value curve. To draft for eliteness, not expected value, earlier picks are more valuable relative to later picks. Futher, as the eliteness cutoff increases, the right tail draft value curve gets steeper and steeper until it nearly matches the fitted trade market curve. In particular, general managers trade draft picks as if they value a draft position by the probability it results in a player whose second contract exceeds $15\%$ of the salary cap. Though we can’t peer into the minds of general managers, if they were trading for eliteness, we would have seen a similar trade market. This is a high bar of eliteness: in our dataset of draft picks from 2013 to 2023, just 16 players ($\approx 0.57\%$ of players) satisfy $Y\geq 15\%$.
Figure 4: Performance value relative to the first pick ($y$-axis) as a function of draft position $x$. The colored lines are right tail probability curves proportional to ${\mathbb{P}}(Y>r|x)$, the dotted black line is proportional to expected performance value ${\mathbb{E}}[Y|x]$, and the solid black line is the fitted trade market curve.
2.2 S curves
[…]
Figure 5: Performance outcome value functions ${g_{\text{step}}}(y)=\mathbb{I}\left(y>r\right)$ (a), two ${g_{\text{s}}}$ curves (b,c), and ${g_{\text{line}}}(y)=y$ (d).
An oddity of the step function ${g_{\text{step}}}$ is that it so sharply characterizes success and failure: $Y=r$ is a total failure and $Y=r+\varepsilon$ is a total success. A weakness of the linear function ${g_{\text{line}}}$ is that it discounts the outsized influence of elite players on winning. For instance, it values the difference between $2\%$ and $1\%$ the same as the difference between $9\%$ and $8\%$. If you believe great players are exponentially more important than typical players for winning games, it is critical to use a nonlinear $g$.
Hence, we propose a nonlinear transformation $g$ between the two extremes of ${g_{\text{step}}}$ and ${g_{\text{line}}}$. To nonlinearly transform performance but in a less sharp way, we use an $s$ curve. We use the CDF of the beta distribution, a sufficiently rich family of $s$ curves, denoted by ${g_{\text{s}}}$ and parameterized by $\alpha$ and $\beta$,
[…]
We visualize two of these $s$ curves in Figures 5b and 5c. For example, ${g_{\text{s}}}$ in Figure 5b assigns near failure to busts and near total success to elite players. The two dashed gray lines indicate where ${g^{\prime}_{\text{s}}}(y)=1$. When $y$ is between the two gray lines, ${g^{\prime}_{\text{s}}}(y)>1$. This reflects, for example, that the difference in value between a good and median player is much larger for ${g_{\text{s}}}$ than for ${g_{\text{line}}}$ (since ${g_{\text{line}}}^{\prime}(y)\equiv 1)$. Similarly, the difference in value between a median and a bad player is much larger for ${g_{\text{s}}}$ than for ${g_{\text{line}}}$.
An $s$ curve ${g_{\text{s}}}$ produces a draft value curve ${v_{\text{s}}}(x)\propto{\mathbb{E}}[{g_{\text{s}}}(Y)|x]$, which we visualize in Figure 6. The $s$ curve draft curves ${v_{\text{s}}}$ (green and purple dashed lines) lie between expected performance (black dotted line) and right tail probability curves (red and blue solid lines). The ${v_{\text{s}}}$ curves lie between the two extremes because they place exponentially increasing value on increasingly good players while still placing nonzero value on less-than-elite production.
Figure 6: The value $v(x)\propto{\mathbb{E}}[g(Y)|x]$ of draft position $x$ relative to the first pick ($y$-axis) for various transformations $g$ of performance (color). Draft curves arising from step function transformations $g={g_{\text{step}}}$ are solid lines, from $s$ curve transformations $g={g_{\text{s}}}$ are dashed lines, and from the linear function $g={g_{\text{line}}}$ are dotted lines.
2.3 Surplus value
Massey and Thaler, (2013), an economics paper at heart, are interested in surplus value, or the difference between first contract performance value and cost. Surplus value accounts for the varying cost of each draft pick (recall the cost curve from Figure 1). Formally, denoting the first contract compensation of draft pick $x$ by ${\mathrm{cost}}(x)$ and its performance outcome by $Y$, surplus value is $S=Y-{\mathrm{cost}}(x)$. Massey and Thaler, (2013) model expected surplus value ${\mathbb{E}}[S|x]$. In the previous sections, we considered nonlinear transformations of performance value that upweighted eliteness. Here, we extend this approach to surplus value, exploring nonlinear transformations of surplus, $v(x)\propto{\mathbb{E}}[g(S)|x]$.
In Figure 7 we visualize expected surplus value (black dashed line) and the probability that surplus value exceeds a threshold (colored lines). Recall the loser’s curse from Massey and Thaler, (2013) and Section 1: according to expected surplus value, last year’s worst teams who pick at the top of the draft have less valuable picks than better teams who pick later in the first round. We visualize this by the large spike in the black dashed line. If you value a pick by the probability it results in a large surplus, the effect size of the loser’s curse diminishes and the surplus peak shifts earlier in the draft.
Figure 7: Surplus value relative to the first pick ($y$-axis) as a function of draft position $x$. The colored lines are right tail probability curves proportional to ${\mathbb{P}}(S>r|x)$ and the dashed black line is proportional to expected surplus value ${\mathbb{E}}[S|x]$.
2.4 Adjusting for position
The performance $Y$ of a player drafted at position $x$ depends on his position, particularly whether he is a quarterback. Quarterbacks command the largest contracts in the NFL and have by far the highest wins above replacement according to PFF WAR (Eager,, 2020). Formally, quarterbacks have a much fatter right tail than other positions (see Figure 19 in Appendix D.2). This is because they have a much higher mean and variance than other positions (see Figure 18 in Appendix D.2). Beyond quarterback, draft curves should also potentially differ across other positions. Previous research suggests it is more difficult to forecast the value of certain positions than other positions. Also, certain positions are inherently more valuable than others (Cole,, 2023; Baldwin,, 2024). Consequently, it is important to adjust for position (${\mathsf{pos}}$), or at least for quarterbacks (${\mathsf{qb}}$), in modeling the density of performance outcomes.
Accordingly, in Appendix D.2 we estimate the conditional density ${\mathbb{P}}(Y|x,{\mathsf{pos}})$ using a Bayesian hierarchical Beta regression model. It is crucial to shrink positional estimates towards a common mean since there are so few datapoints for each position. The standard errors of the positional conditional densities (and conditional means) are huge and need to be accounted for. We don’t find large or significant differences across non-quarterback positions (see Figure 18). But, quarterbacks have significantly higher conditional mean and variance curves (and hence fatter right tails) than non-quarterbacks. Thus, in formulating draft trade value curves, we consider just quarterbacks and non-quarterbacks. We define the non-quarterback conditional density by averaging over all the non-quarterback positions, ${\mathbb{P}}(Y|x,{\mathsf{not\ qb}}):=\frac{1}{\#{\mathsf{pos}}-1}\sum_{k\neq{\mathsf{qb}}}{\mathbb{P}}(Y|x,{\mathsf{pos}}=k)$.
In Figure 8 we visualize performance value curves for quarterbacks and non-quarterbacks. We see the same general trend from Figure 4: higher eliteness cutoffs correpond to steeper curves. Further, the higher the cutoff of eliteness, the less a non-${\mathsf{qb}}$ pick is worth relative to a first pick ${\mathsf{qb}}$. For instance, the probability that the second contract of a non-${\mathsf{qb}}$ drafted first exceeds $15\%$ of the cap is about $10\%$ of that of a ${\mathsf{qb}}$ (see the purple lines). In contrast, the expected performance value of a non-${\mathsf{qb}}$ drafted first is about $60\%$ of that of a ${\mathsf{qb}}$ (see the dashed black lines). This makes sense: a non-${\mathsf{qb}}$ drafted first will on average provide less value than a ${\mathsf{qb}}$ drafted first, but nowhere near as much less value than if we value performance by eliteness. For large cutoffs of eliteness, the value of later non-${\mathsf{qb}}$ picks quickly converges to $0\%$ of a first pick ${\mathsf{qb}}$. This makes sense: just about the only way to get an extremely elite, far-right tail performance outcome is from a ${\mathsf{qb}}$.
Figure 8: Performance value relative to the first ${\mathsf{qb}}$ pick ($y$-axis) as a function of draft position $x$ and position (facet). The colored lines are right tail probability curves proportional to ${\mathbb{P}}(Y>r|x,{\mathsf{pos}})$ and the dashed black line is proportional to expected performance value ${\mathbb{E}}[Y|x,{\mathsf{pos}}]$.
In Figure 9 we visualize surplus value curves for quarterbacks and non-quarterbacks. According to expected surplus value, there is a loser’s curse for ${\mathsf{qb}}$s and non-${\mathsf{qb}}$s, as seen by the hump in the dashed black line: last year’s worst teams who pick at the top of the draft have less valuable picks than better teams who pick later in the first round. If you are interested in an elite ${\mathsf{qb}}$, however, there is no loser’s curse: the surplus peak all but disappears for right tail probability surplus curves. The far-right tails of performance $Y$ are so fat for top pick ${\mathsf{qb}}$’s that even the high cost of top picks doesn’t massively diminish their value.333 See Figure 20 in Appendix D.2 to visualize the right tails of the conditional surplus densities for ${\mathsf{qb}}$s and ${\mathsf{not\ qb}}$s. For non-quarterbacks, the loser’s curse remains, even if you are interested in eliteness. Surplus value peaks in the middle of the first round and decays across the draft. This makes sense: early picks’ compensation are priced for the ultra high mean and variance of quarterbacks. On average, this price becomes too high for other positions.
Figure 9: Surplus value relative to the first ${\mathsf{qb}}$ pick ($y$-axis) as a function of draft position $x$ and position (facet). The colored lines are right tail probability curves proportional to ${\mathbb{P}}(S>r|x,{\mathsf{pos}})$ and the dashed black line is proportional to expected surplus value ${\mathbb{E}}[S|x,{\mathsf{pos}}]$.
3 Discussion
Traditional NFL draft value curves are expected value curves: the value of a draft position is proportional to expected performance or surplus value. Massey and Thaler, (2013), who first constructed these curves, had two main findings. First, they coined the loser’s curse: top draft picks, belonging to last year’s worst teams, produce less surplus value on average than draft picks later in the first round that belong to better teams. Second, the draft value curve implied by the trade market is much steeper than expected performance and surplus value curves. Thus, according to those expected value curves, general managers trade suboptimally: they don’t trade down often enough, and they often overpay when they trade up.
Massey and Thaler’s analysis from over a decade ago completely replicates to recent data (2013-2023). Even though their work has been public and widely known for a long time, general managers’ draft behavior has remained largely unchanged. Why? It may be that general managers are stubborn and not wise. Alternatively, there is a sound economic reason why average future player value (at a draft position) doesn’t determine trade value. We are partial to the latter explanation.
The goal is to win a Super Bowl, or to be the best team of 32 teams. To win a Super Bowl, we suspect that elite players, and particularly elite quarterbacks, are exponentially more valuable than average players. Thus, we construct draft value curves that value eliteness. The most intuitive example is right tail probability: the value of a draft position is proportional to the probability that the performance or surplus of a drafted player exceeds a threshold. We find that a steep right tail probability curve nearly matches the curve implied by the trade market, suggesting that general managers may want elite players, not average players. They may be willing to pay quite a lot for draft picks that are likely to be elite NFL performers. Then, we adjust for position, creating separate expected value and right tail probability draft curves for quarterbacks and non-quarterbacks. We find that, if you are looking for an elite quarterback, there is no loser’s curse. If you are looking for elite talent, picks from lower draft picks aren’t worth much and the high cost of a top draft pick is justified.
The right objective function is Super Bowl win probability: the value of a draft position should be proportional to the probability that a drafted player increases the probability you win the Super Bowl in the next $t$ years. We do not solve this problem because it is extremely difficult. How the addition of a draft pick changes a team’s Super Bowl win probability depends on the current composition of his team (in particular, projected current and future talent level at each position), injury risk, salary cap considerations, and projected outcomes of players in the draft. This results in an extremely fascinating and complicated problem, which we leave to future work.
The curves detailed in this work are simple proxies of Super Bowl win probability. Which proxy is right to use depends on the aforementioned factors (team composition, salary cap, and projected outcomes of players in the draft) and a team’s goals. In general, teams should trade draft picks based on some combination of eliteness and expected value. For example, consider the general manager of a below average team without an elite quarterback (think the 2023 Bears) whose goal is to win the Super Bowl sometime soon. Intuitively, drafting a median or even a good player barely increases the probability of achieving his goal, but drafting a potentially highly elite player (such as Caleb Williams) significantly increases it. Thus, we recommend using a steep trade value chart. This may explain why the Bears did not relinquish its number one pick in the 2024 draft. Now, consider the general manager of a consecutive Super Bowl winning team featuring an elite quarterback, tight end, and defensive tackle (think the 2023 Chiefs) whose goal is to win the Super Bowl sometime soon. We recommend using a shallower curve than that of the Bears. The difference in Super Bowl win probability between adding an average player and a bust is larger for the Chiefs than for the Bears. An average player can provide a real impact to the Chiefs winning another Super Bowl.
Though our analysis builds upon the draft curve literature, it is not without limitations, and there are many avenues for future work. First, the methodology underlying the fitted trade market curve (the Weibull curve from Appendix C) has imperfections, which we detail in the Appendix. In particular, if the team trading down uses a different value function than the team trading up, how do we explain trading? Second, in future work, we suggest exploring curves built from the sum of performance and surplus, rather than each one individually. When you draft a player, you benefit from the value of their performance in addition to the extra cash from a cheap rookie contract. Third, a major assumption employed in this work and in previous work is additivity: the value of a bundle of draft picks $v(\{x_{j}\})$ equals its sum, $\sum_{j}v(x_{j})$. Concretely, in terms of right tail probability, we consider the value of a bundle of picks by $\sum_{j}{\mathbb{P}}(Y_{j}>r)$, though the relationship between the picks may be non-additive or something more general, like ${\mathbb{P}}(f(Y_{1},...,Y_{n})>r)$. We leave an exploration of non-additve draft curves to future work. Fourth, in this study we consider just first contract performance and surplus value. First contract value is an incomplete measure of the total value he provides to his team, as it excludes the rest of his career. Finally, we did not address a selection bias that was explored in Eager, (2024): worse teams draft players earlier in each round of the draft than better teams, and a team’s quality may influence the ultimate performance outcome of a drafted player. A more elaborate construction of NFL draft trade value curves should account for this selection bias, which we leave to future work.
Acknowledgments
The authors thank Cade Massey, Christopher N. Avery, and Blake Zilberman for giving the authors a reason to explore draft value curves. We also thank Cade and Chris for sending NFL draft data for an early version of this project.
SUPPLEMENTARY MATERIAL
Appendix A Our code
Our analysis in this paper is reproducible via the code that is publicly available on Github at https://github.com/snoopryan123/NFL_draft_chart_Ryan.
Appendix B Measuring a player’s performance value
Each play in American football consists of a complex interplay between 22 players of varying roles interacting in continuous time and space, making it difficult to measure how a player’s individual performance impacts the ultimate outcome of the play. For instance, a linebacker may impact the outcome of a play simply by occupying space, causing the ball carrier to move in a different direction, and there is no easy way to measure the impact of that action. Since there is no easy or industry-standard way of measuring player performance on a given play, particularly for off-the-ball players, analysts use simpler proxies of performance value. Importantly, these proxies measure the performance of players of all positions on the same scale. Each of these are imperfect measures with their own drawbacks.
In free agency, NFL players who are not bound by a contract can sign a new contract with any team. Assuming the free agent market is efficient, it is a reflection of the true performance value of players. Hence, one popular measure of performance value is inflation-adjusted salary. Baldwin, (2024) adjusts for inflation by measuring salary as a percentage of the salary cap, referred to as APY cap percentage (APY means average salary per year). Fitzgerald and Spielberger, (2024) measure a player’s performance value as his salary as a percentage of the average of the top five salaries at his position. For example, in 2024 the top five quarterback salaries have an APY of \$51.2 million, so they grade Mahomes’ APY of \$45 million as an 88%. A drawback of measuring a player’s value relative to other players at his position is it doesn’t account for positional value, or the fact that some positions are inherently more valuable than others (e.g., an average quarterback is much more valuable than an average running back). A drawback of using the free agent market to measure performance is that the market is a reflection of projected future value, not past performance value. A running back who performs very well in his first contract may not receive a large second contract because running backs are injury prone and age quickly. For example, Saquon Barkley, one of the best running backs in the NFL during his rookie contract, signed a modest second contract in 2023 worth \$10.1 million for one year. His free agent contract surely understates his first contract performance value.
To measure past performance value rather than projected future performance value, Massey and Thaler, (2013) use the number of games a player started and whether he made the Pro Bowl. They then map this performance to salary by regressing a player’s second contract compensation (the value of a player’s first free agent contract) on his number of starts and Pro Bowls in his first contract (see Section 6.2 of their paper for the particulars of this regression). A drawback of their approach is that the number of starts and Pro Bowls is a crude measure of player performance. Two linebackers can have the same number of starts and Pro Bowls but have vastly different impacts on the games they played, for instance by having a different number of tackles or yards allowed.
Stuart, (2012) measures player performance using Pro Football Reference’s Approximate Value (AV) (Pro Football Reference,, 2024). AV maps more granular position-specific performance measures to a common scale. For offensive players, AV apportions the points per drive scored by a team’s offense to each of its offensive players in proportion to position-specific performance measures. For skill position players, AV uses yards to measure performance (e.g., rushing yards for running backs and pasing yards for quarterbacks). AV uses more crude measures for offensive lineman (e.g., games played, games started, and All Pro). AV divides credit for an offensive team’s points scored per drive using ad hoc multipliers (e.g., the offensive line gets $5/11$ of the offensive credit since there are usually 5 offensive lineman on the field). AV works analogously for defensive players. Drawbacks of AV are that it is too simple, too ad hoc, and uses crude performance measures for non-skill position players.
Pro Football Focus, (2024) (known as PFF) measures player performance using Wins Above Replacement (WAR) (Eager,, 2020). WAR estimates an individual player’s contribution to team performance on the scale of team wins relative to a replacement-level player (e.g., see Brill and Wyner, (2024) for an exposition of WAR for starting pitchers in baseball). PFF grades the performance of every player in every football game from -2 to +2 in increments of $1/2$, where 0 represents average performance. They then estimate WAR from these PFF grades using a Massey rating system (Massey,, 1997). A drawback of PFF WAR is that it is a model-based metric and it is unclear how good or reliable this model is at measuring player performance. Some people take issue with PFF grades, which are subjective measures of player performance made by humans who watch the tape. For instance, a Pro Bowl NFL offensive lineman once told us that PFF often grades him incorrectly because it doesn’t know the job his coach told him to do. He should’ve received a perfect grade for doing his job, which is not publicly available knowledge, even though watching the tape may make it seem like he played suboptimally.
Appendix C Estimating the trade market value of draft picks
We follow the approach from Massey and Thaler, (2013) to estimate the relative value of draft picks that is implied by the NFL trade market. This trade market value function ${v_{\text{market}}}$ is a function of a bundle of picks $\{x_{j}\}_{j=1}^{n}$, where each pick $x_{j}$ belongs to $\{1,...,256\}$. Without loss of generality, assume that picks in a bundle are ordered from earliest to latest pick, $x_{1}\leq...\leq x_{n}$. Also, ${v_{\text{market}}}(\{1\})={v_{\text{market}}}(1)=1$ so that value is relative to the first pick.
Massey and Thaler, (2013) make simplifying assumptions in fitting ${v_{\text{market}}}$. First, they assume additivity,
${v_{\text{market}}}(\{x_{j}\}_{j=1}^{n})=\sum_{j=1}^{n}{v_{\text{market}}}(x_{j}).$ (C.1)
Second, they suppose that all teams use the same market value function ${v_{\text{market}}}$ to value draft picks. Third, they assume that trades in the market are fair, so the market value of both teams’ bundles of picks are equal. Formally, denoting the bundle of picks initially belonging to the team trading down by $\{x_{j}^{(d)}\}_{j=1}^{n_{d}}$ and that of the team trading up by $\{x_{j}^{(u)}\}_{j=1}^{n_{u}}$, the trade is fair if
$\sum_{j=1}^{n_{d}}{v_{\text{market}}}(\{x_{j}^{(d)}\})=\sum_{j=1}^{n_{u}}{v_{\text{market}}}(\{x_{j}^{(u)}\}).$ (C.2)
Finally, they assume the market value of picks follows a Weibull distribution,
${v_{\text{market}}}(x)=e^{-\lambda(x-1)^{\beta}},$ (C.3)
where $\lambda$ and $\beta$ are parameters to be estimated. They also consider a Weibull model with a discount rate $\rho$ for a pick $N$ years into the future,
${v_{\text{market}}}(x,N)=e^{-\lambda(x-1)^{\beta}}\cdot(1+\rho)^{N}.$ (C.4)
Combining equations (C.2), (C.1), and (C.3), they calculate
$\log(x^{(d)}_{1})=\log\bigg{(}\bigg{[}-\frac{1}{\lambda}\log\bigg{(}\sum_{j=1}^{n_{u}}e^{-\lambda(x_{j}^{(u)}-1)^{\beta}}-\sum_{j=2}^{n_{d}}e^{-\lambda(x_{j}^{(d)}-1)^{\beta}}\bigg{)}\bigg{]}^{1/\beta}\bigg{)},$ (C.5)
which expresses the value of the top pick in the trade acquired by the team trading down in terms of the other picks in the trade. They estimate the parameters $\lambda$ and $\beta$ in Equation (C.5) using nonlinear regression. They use a similar approach to find a trade market value curve that incorporates the discount rate.
The aforementioned assumptions range from quite strong to very strong and deserve to be examined. The fourth assumption, that ${v_{\text{market}}}$ follows a Weibull distribution, seems to be the most mild, as the Weibull is a sufficiently expressive family that can capture many different plausible draft value curves. The first assumption, additivity, is wrong but mild. For instance, the value of a bundle of $12$ picks is not additive because not all of these players can play on the field at the same time. But additivity doesn’t seem to be a terrible assumption because most trades involve just a few picks. The third assumption is wrong; they probably meant to assume that observed trades are fair on average:
$\sum_{j=1}^{n_{d}}{v_{\text{market}}}(\{x_{j}^{(d)}\})=\sum_{j=1}^{n_{u}}{v_{\text{market}}}(\{x_{j}^{(u)}\})+\varepsilon,$ (C.6)
where $\varepsilon$ is mean zero noise. The analogue of Equation C.5 under this new assumption is
$\log(x^{(d)}_{1})=\log\bigg{(}\bigg{[}-\frac{1}{\lambda}\log\bigg{(}\sum_{j=1}^{n_{u}}e^{-\lambda(x_{j}^{(u)}-1)^{\beta}}-\sum_{j=2}^{n_{d}}e^{-\lambda(x_{j}^{(d)}-1)^{\beta}}+\varepsilon\bigg{)}\bigg{]}^{1/\beta}\bigg{)},$ (C.7)
which is not equal to
$\log(x^{(d)}_{1})=\log\bigg{(}\bigg{[}-\frac{1}{\lambda}\log\bigg{(}\sum_{j=1}^{n_{u}}e^{-\lambda(x_{j}^{(u)}-1)^{\beta}}-\sum_{j=2}^{n_{d}}e^{-\lambda(x_{j}^{(d)}-1)^{\beta}}\bigg{)}\bigg{]}^{1/\beta}\bigg{)}+\varepsilon.$ (C.8)
The nonlinear regression they employ implicitly uses the form from Equation (C.8) even though they should fit from Equation (C.7). Notably, however, the second assumption (that all teams use the same market value function ${v_{\text{market}}}$ to value draft picks) is the most egregious. Each general manager has his own value function, which changes from year to year and can be skewed by meddling owners or coaches. We could potentially treat the value function employed in each trade as a draw centered at ${v_{\text{market}}}$, with multiple levels of variance for each general manager and each year. There is likely not enough data to estimate such a hierarchical model well and its standard error would be large.
Despite these issues with the simplifying assumptions, we replicate the analysis discussed above from a dataset consisting of all draft trades consisting purely of draft picks from 2006-2023. In light of the aforementioned flaws, we use the fitted ${v_{\text{market}}}$ curve to tell a general story about the steepness of the trade market. We take it with a grain of salt without obssessing over the exact shape of the curve or its minutae. We display the ${v_{\text{market}}}$ from Massey and Thaler, (2013) and from our replication, with and without the discount factor, in Figure 10.
Figure 10: Weibull trade market value curves ${v_{\text{market}}}$ according to Massey and Thaler, (2013) and our replication.
Appendix D Estimating the conditional density of the outcome of a draft pick
D.1 Position-agnostic conditional density
In Figure 11 we visualize the empirical conditional density of $Y$ given $x$. The empirical densities are extremely wobbly and noisy due to limited data for each individual draft pick. But there is a general trend: we see a spike near zero that shifts upwards as draft pick increases and a right tail that shifts leftwards and gets smaller as draft pick increases. This motivates our “Bayesian spike plus Beta regression” model of ${\mathbb{P}}(Y|x)$.444 We initially tried a lone Beta regression model without the spike near zero, which less successfully fit the right tail.
Figure 11: Empirical distribution (histogram and smoothed density) of performance outcome $Y$ (second contract APY cap percentage, $x$-axis) for each draft pick $x$ (facet).
We model the spike near zero, the bust spike, using logistic regression. We let a bust be a player whose performance outcome $Y$ is smaller than some cutoff ${y_{\text{bust}}}$. We found that ${y_{\text{bust}}}=0.005$ (or $0.5\%$ of the cap) works well. We model the bust probability ${\mathrm{bp}}(x):={\mathbb{P}}(Y\leq{y_{\text{bust}}}|x)$ by
$Y\leq{y_{\text{bust}}}|x\sim{\mathrm{Bernoulli}}({\mathrm{bp}}(x)),$ (D.1)
where
${\mathrm{bp}}(x)={\mathrm{logistic}}(\alpha_{0}+\alpha_{1}\cdot x).$ (D.2)
Next, we model the right tail using Beta regression. We view the Beta distribution as a natural way to model $Y$ because it lies in $[0,1]$ and it seems reasonable when we eyeball the empirical densities (á la Figure 11). We model
$Y|x,Y>{y_{\text{bust}}}\sim{\mathrm{Beta}}(\mu(x),\phi(x)).$ (D.3)
Here, we use the mean-precision parameterization of the Beta distribution. The mean $\mu$ and precision $\phi$ are related to the traditional Beta distribution parameters $\mathrm{shape}_{1}$ and $\mathrm{shape}_{2}$ by $\mathrm{shape}_{1}=\mu\phi$ and $\mathrm{shape}_{2}=(1-\mu)\phi$. We let the mean and precision parameters vary as draft pick $x$ varies via $\mu(x)$ and $\phi(x)$ to capture that different draft picks $x$ yield different outcome distributions. The conditional mean is ${\mathbb{E}}[Y|x,Y>{y_{\text{bust}}}]=\mu(x)$ and the conditional variance is $\mathrm{Var}[Y|x,Y>{y_{\text{bust}}}]=\mu(x)(1-\mu(x))/(1+\phi(x))$. So, the precision is inversely proportional to the variance. We use spline regression to model the conditional mean,
$\mu(x)={\mathrm{logistic}}(\widetilde{x}^{\top}\beta),$ (D.4)
where $\widetilde{x}$ is a B-spline basis.555 The ${\mathsf{R}}$ code we used to generate the spline basis is splines::bs($x$, degree=3, df=4, intercept=T). We model the conditional precision by
$\phi(x)=\exp(\gamma_{0}+\gamma_{1}\cdot x).$ (D.5)
To capture uncertainty in the parameters, we use a Bayesian model, placing priors on the parameters. We put a weakly informative Normal prior with mean 0 and s.d. 10 on each of the $\alpha$, $\beta$, and $\gamma$ parameters. Finally, we assume that $Y$ is uniformly distributed within the bust spike in a way that doesn’t vary with $x$, $Y|x,Y\leq{y_{\text{bust}}}\sim{\mathrm{Unif}}[0,{y_{\text{bust}}}]$.666 We assess the reasonableness of this assumption later in this section. But this assumption is benign because when $Y\leq{y_{\text{bust}}}$, $Y$ is essentially 0 and so negligbly impacts the right tail probability calculations that form the draft value curves that we detail in this paper.
Because the posterior distribution of $(\alpha,\beta,\gamma)$ is not analytically tractable, we use Markov Chain Monte Carlo (MCMC) to draw approximate samples from the posterior distribution. We implement our sampler in Stan (Carpenter et al.,, 2017) and perform our MCMC simulation using the rstan package (Stan Development Team,, 2022). To obtain our posterior samples, we run one MCMC chain for 2,500 iterations. After discarding the first 1,250 iterations of each chain as “burn-in”, the Gelman-Rubin $\hat{R}$ statistic for each parameter is less than $1.1$, suggesting convergence (Gelman and Rubin,, 1992).
In Figure 12 we see that our estimates of the conditional mean $\mu(x)$, conditional standard deviation ${\mathrm{sd}}(x)$, and bust probability ${\mathrm{bp}}(x)$ fit the data well. In Figure 13 we see that our estimates of the right tail ${\mathbb{P}}(Y|x,Y>{y_{\text{bust}}})$ fit the data well enough. As expected, the right tail shifts leftwards as draft pick increases. In Figure 14 we visualize the conditional density ${\mathbb{P}}(Y|x)$ for values $Y$ in the right tail $Y>{y_{\text{bust}}}$. By Bayes rule, this is equal to ${\mathbb{P}}(Y|x,Y>{y_{\text{bust}}})\cdot(1-{\mathrm{bp}}(x))$. We see that the right tail, measuring the probability of drafting an elite player, shifts leftwards and decreases in size as draft pick increases. The posterior credible intervals of these density curves are fairly narrow because the information of our probability model is distilled into just a few parameters. In Figure 15 we visualize the conditional density of surplus $S=Y-{\mathrm{cost}}$. The high cost of the first few picks shifts those curves to the left relative to curves of other picks.
Figure 12: Left: our estimate of the conditional mean $x\mapsto\mu(x)={\mathbb{E}}[Y|x,Y>{y_{\text{bust}}}]$ given not a bust. Middle: our estimate of the conditional standard deviation $x\mapsto{\mathrm{sd}}(x)={\mathrm{sd}}(Y|x,Y>{y_{\text{bust}}})$ given not a bust. Right: our estimate of bust probability $x\mapsto{\mathrm{bp}}(x)={\mathbb{P}}(Y\leq{y_{\text{bust}}}|x)$. The solid black line is the posterior mean, the gray shaded region is the $95\%$ credible interval, and the black dots are the empirical values for each draft pick $x$.
Figure 13: The solid black lines are the fitted conditional density ${\mathbb{P}}(Y|x,Y>{y_{\text{bust}}})$ (the posterior mean of our model). This overlays the empirical distribution (gray bars and gray lines) for each draft pick.
Figure 14: For values $Y$ in the right tail $Y>{y_{\text{bust}}}$, the estimated conditional density ${\mathbb{P}}(Y|x)={\mathbb{P}}(Y|x,Y>{y_{\text{bust}}})\cdot(1-{\mathrm{bp}}(x))$ for various values of draft pick $x$ (color). Figure (a) displays the posterior mean density curves and (b) displays 95% credible intervals of these curves.
Figure 15: For values $Y$ in the right tail $Y>{y_{\text{bust}}}$, the estimated conditional surplus density ${\mathbb{P}}(S|x)={\mathbb{P}}(Y-{\mathrm{cost}}(x)|x,Y>{y_{\text{bust}}})\cdot(1-{\mathrm{bp}}(x))$ for various values of draft pick $x$ (color) in the first round (a) and throughout the draft (b).
Finally, in Figure 16 we assess the assumption that the bust spike conditional distribution of $Y|x,Y\leq{y_{\text{bust}}}$ is uniform in a way that doesn’t vary with $x$. The empirical conditional mean stays essentially constant over $x$ and the empirical conditional variance slightly decreases linearly in $x$. This admittedly renders the uniform assumption imperfect. Nonetheless, we ultimately feel it is good enough. The assumption is benign because when $Y\leq{y_{\text{bust}}}$, $Y$ is essentially 0 and so negligbly impacts the right tail probability calculations that form the draft trade value curves that we detail in this paper.
Figure 16: The empirical mean (left) and standard deviation (right) of $Y$ for each draft pick $x$ conditional on being a bust $Y\leq{y_{\text{bust}}}$.
D.2 Position-specific conditional density
In this section, we model the conditional density of performance outcome $Y$ given draft pick $x$ and position ${\mathsf{pos}}$. We extend the “Bayesian spike plus Beta regression” model from Appendix D.1 to adjust for position. We use 11 position groups: quarterback (QB), wide receiver (WR), running back / fullback (RB/FB), tight end (TE), offensive tackle (OT), interior offensive lineman (IOL), interior defensive lineman (DL), edge (ED), linebacker (LB), cornerback (CB), and safety (S). We filter out all draft picks from our dataset belonging to other positions (e.g., punter, kicker, and long snapper).
It is imperative to use a Bayesian model to adjust for position, rather than fitting separate curves for each position, since there is such little data within each position group. Across each of the position groups, the median sample size in our dataset is 269 players, with a maximum of 350 and a minimum of 127. Spread across 256 total draft pick positions, this is not an abundance of data. A Bayesian hierarchical model is an excellent choice here because it allows us to capture uncertainty in position curves while also shrinking towards an overall mean.
We devise a Bayesian hierarchical model that includes position-specific parameters for each of the $\alpha$, $\beta$, and $\gamma$ parameters from the previous model. We model the right tail conditional density by
$Y|x,{\mathsf{pos}},Y>{y_{\text{bust}}}\sim{\mathrm{Beta}}(\mu(x,{\mathsf{pos}}),\phi(x,{\mathsf{pos}})),$ (D.6)
where
$\mu(x,{\mathsf{pos}})={\mathrm{logistic}}(\widetilde{x}^{\top}\beta_{\mathsf{pos}})$ (D.7)
and
$\phi(x,{\mathsf{pos}})=\exp(\gamma_{0,{\mathsf{pos}}}+\gamma_{1,{\mathsf{pos}}}\cdot x).$ (D.8)
As before, $\widetilde{x}$ is a B-spline basis on the $x$ variable. We model bust probability by
${\mathrm{bp}}(x,{\mathsf{pos}}):={\mathbb{P}}(Y\leq{y_{\text{bust}}}|x,{\mathsf{pos}})={\mathrm{logistic}}(\alpha_{0,{\mathsf{pos}}}+\alpha_{1,{\mathsf{pos}}}\cdot x).$ (D.9)
We shrink the position-specific coefficients $\alpha_{\ast,{\mathsf{pos}}}$, $\beta_{\ast,{\mathsf{pos}}}$, and $\gamma_{\ast,{\mathsf{pos}}}$ towards overall mean parameters $\alpha_{\ast}$, $\beta_{\ast}$, and $\gamma_{\ast}$. For example, we model the position-specific coefficients $\alpha_{0,{\mathsf{pos}}}$ as a draw from a Normally distributed prior with mean $\alpha_{0}$ and standard deviation $\tau_{\alpha_{0}}$. We put a weakly informative Normal prior with mean 0 and s.d. 10 on the prior mean parameter $\alpha_{0}$. We put a mild positive standard Normal prior on the prior standard deviation parameter $\tau_{\alpha_{0}}$. The priors and hyperpriors are similar for the other parameters.
We again use Stan to draw approximate posterior samples of the parameters $(\alpha,\beta,\gamma)$, running one chain for 50,000 iterations. After discarding the first 25,000 iterations of each chain as “burn-in”, the Gelman-Rubin $\hat{R}$ statistic for each parameter is less than $1.1$, suggesting convergence (Gelman and Rubin,, 1992).
In Figure 17a we see that our estimate of the conditional mean $\mu(x,{\mathsf{pos}})$ fits the data well. The empirical standard deviation values in Figure 17b we are much more sparse than the empirical conditional mean values in Figure 17a. This is because, for each position ${\mathsf{pos}}$ and for many draft picks $x$, just one player of position ${\mathsf{pos}}$ was drafted at $x$. Thus, there is no empirical standard deviation at many combination of $x$ and ${\mathsf{pos}}$. Nonetheless, based on the data we have observed, our estimate of the conditional standard deviation ${\mathrm{sd}}(x,{\mathsf{pos}})$ fits the data reasonably well.
Figure 17: Figure (a): our estimate of the conditional mean $\mu(x,{\mathsf{pos}})={\mathbb{E}}[Y|x,{\mathsf{pos}},Y>{y_{\text{bust}}}]$ given not a bust. Figure (b): our estimate of the conditional standard deviation ${\mathrm{sd}}(x,{\mathsf{pos}})={\mathrm{sd}}(Y|x,{\mathsf{pos}},Y>{y_{\text{bust}}})$ given not a bust. The solid black line is the posterior mean, the gray shaded region is the $95\%$ credible interval, and the black dots are the empirical values for each draft pick $x$.
In Figure 18a we see that the posterior means of the conditional mean, standard deviation, and bust probability curves vary by position. WR and OT have the highest non-quarterback conditional mean lines and RB/FB, S, LB, and TE have the lowest non-quarterback conditional mean lines. WR and ED have the highest non-quarterback conditional s.d. lines and RB/FB, S, and TE have the lowest non-quarterback conditional s.d. lines. The bust probability lines are clumped into two groups, one about $0.05$ probability higher than the other. In Figure 18b we add 95% credible intervals (the shaded regions) to these curves. The credible intervals are wide due to a dearth of data across the draft at each individual position. The non-quarterback mean and s.d. curves jumble into one big clump. Closer inspection may reveal two clumps of non-quarterback positions, one of whose curves lie slightly above the other’s, but the effect size of this difference is small. Thus, in Figure 18c we average all of the non-quarterback curves into the red curves, $\mu(x,{\mathsf{not\ qb}})=\frac{1}{\#{\mathsf{pos}}-1}\sum_{{\mathsf{pos}}\ k\neq{\mathsf{qb}}}\mu(x,{\mathsf{pos}})$ and similarly for ${\mathrm{sd}}(x,{\mathsf{not\ qb}})$ and ${\mathrm{bp}}(x,{\mathsf{not\ qb}})$. As expected, quarterbacks have a higher mean than other positions. What’s more interesting is that quarterbacks have astoundingly higher variance than other positions. This is what drives the fat far right tail of top pick quarterbacks.
Figure 18: Figure (a): The posterior mean estimates (lines) of $\mu(x,{\mathsf{pos}})$, ${\mathrm{sd}}(x,{\mathsf{pos}})$, and ${\mathrm{bp}}(x,{\mathsf{pos}})$ as a function of draft pick $x$ and position ${\mathsf{pos}}$ (color). In Figure (b) we add 95% credible intervals (the shaded regions) to these curves. Figure (c): These curves for ${\mathsf{pos}}\in\{{\mathsf{qb}},{\mathsf{not\ qb}}\}$.
In Figure 19 we visualize the conditional density ${\mathbb{P}}(Y|x,{\mathsf{pos}})$ for ${\mathsf{pos}}\in\{{\mathsf{qb}},{\mathsf{not\ qb}}\}$ (where the ${\mathsf{not\ qb}}$ density is the average of all the non-quarterback positional densities). As the draft progresses (Figure 19b), for both ${\mathsf{qb}}$’s and ${\mathsf{not\ qb}}$’s the right tail shifts leftwards and decreases in size as draft pick increases. The ${\mathsf{qb}}$ right tails are fatter than those for ${\mathsf{not\ qb}}$. For picks in the first round (Figure 19a), the ${\mathsf{qb}}$ densities have an incredibly fat right tail, significantly fatter than those of ${\mathsf{not\ qb}}$. This reflects the immense value of top pick quarterbacks: these top picks yield the highest probability of acquiring the most elite players in the game, who provide an outsize influence on winning the Super Bowl. In Figure 20 we visualize the conditional density of surplus $S=Y-{\mathrm{cost}}$ for ${\mathsf{qb}}$ and ${\mathsf{not\ qb}}$. For non-quarterbacks, the high cost of the first few picks shifts the right tails of ${\mathsf{not\ qb}}$ to the left of those of later picks, creating a non-decreasing surplus value curve even with our nonlinear performance outcome value functions (like right tail probability). For quarterbacks, even after accounting for cost the right tails of top picks remain fatter than those of later picks. This reflects the immense value of top pick ${\mathsf{qb}}$’s.
Figure 19: For values $Y$ in the right tail $Y>{y_{\text{bust}}}$, the estimated conditional density ${\mathbb{P}}(Y|x,{\mathsf{pos}})={\mathbb{P}}(Y|x,{\mathsf{pos}},Y>{y_{\text{bust}}})\cdot(1-{\mathrm{bp}}(x,{\mathsf{pos}}))$ for ${\mathsf{pos}}\in\{{\mathsf{qb}},{\mathsf{not\ qb}}\}$ (facet) and various values of draft pick $x$ (color) in the first round (a) and throughout the draft (b).
Figure 20: For values $Y$ in the right tail $Y>{y_{\text{bust}}}$, the estimated conditional surplus density ${\mathbb{P}}(S|x,{\mathsf{pos}})={\mathbb{P}}(Y-{\mathrm{cost}}(x)|x,{\mathsf{pos}},Y>{y_{\text{bust}}})\cdot(1-{\mathrm{bp}}(x,{\mathsf{pos}}))$ for ${\mathsf{pos}}\in\{{\mathsf{qb}},{\mathsf{not\ qb}}\}$ (facet) and various values of draft pick $x$ (color) in the first round (a) and throughout the draft (b).